# Exact output tracking for the one-dimensional heat equation and applications to the interpolation problem in Gevrey classes of order 2

### v1

@@ A.2 A proof of Mitjagin’s Theorem
where, as recalled, $(f_{n})$ are the coefficients of the entire function $\Phi$ built in Lemma 28. However, for a sequence $(a_{n})$ of complex numbers and $\phi\in C^{\infty}_{\beta B}$, we are only able to make the series $\sum a_{n}\phi^{(n)}(0)$ convergent under the assumption
@@
By the Cauchy-Schwarz inequality, we deduce that the integral (56) thus converges, that (56) defines a functional on $L^{2}_{\beta B}$, and the first estimate in (57) follows.
@@
Similarly to in the proof of Theorem 27, we deduce that
@@
Denote $X=L^{2}_{\beta,\Gamma_{\beta}B-\epsilon/2}$, from (60) and the Douglas’s Lemma we have
@@
From Proposition 30, the entire function $\Phi$ constructed in Lemma 28 belongs to $L^{2}(\mathbb{R},\omega^{-2}(\xi)d\xi)$, and it has order $\leq 1/\beta$ on the complex plane. From [Borichev, Theorem 3.1], there exists a sequence $(P_{j})$ of polynomials such that $P_{j}\rightarrow\Phi$ for the $L^{2}(\mathbb{R},\omega^{-2}(\xi)d\xi)$-norm. For every $j$ we put
@@ Appendix B Proof of the Lemma on the discrete version of the Laplace method
where the $O$ term is uniform in $n\geq 1$ and $\delta>0$ small enough. Thus, for $\delta>0$ be fixed and small enough, we have
@@
We are thus left to estimate the remaining two series on the right-hand side of (70). Until the end of the proof, $\kappa$ is any positive number which depends on $\delta$ and $x_{0}$, changing line to line, small enough so that the written asymptotics are valid. To bound above the first series we use a comparison between series and integrals:
@@
and we conclude, similarly to the previous step, that

### v2

@@ A.2 A proof of Mitjagin’s Theorem
where $(f_{n})$ are the coefficients of the entire function $\Phi$ built in Lemma 31. However, for a sequence $(a_{n})$ of complex numbers and $\phi\in C^{\infty}_{\beta B}$, we are only able to make the series $\sum a_{n}\phi^{(n)}(0)$ convergent under the assumption
@@
From the Cauchy-Schwarz inequality, we deduce that the integral (58) thus converges, that (58) defines a functional on $L^{2}_{\beta B}$, and the first estimate in (59) follows.
@@
Similarly as in the proof of Theorem 30, we deduce that
@@
Denote $X=L^{2}_{\beta,\Gamma_{\beta}B-\epsilon/2}$, from (62) and Douglas’s Lemma we have
@@
From Proposition 33, the entire function $\Phi$ constructed in Lemma 31 belongs to $L^{2}(\mathbb{R},\omega^{-2}(\xi)d\xi)$, and it has order $\leq 1/\beta$ on the complex plane. From [Borichev, Theorem 3.1], there exists a sequence $(P_{j})$ of polynomials such that $P_{j}\rightarrow\Phi$ for the $L^{2}(\mathbb{R},\omega^{-2}(\xi)d\xi)$-norm. For every $j$, we put
@@ Appendix B Proof of the Lemma on the discrete version of the Laplace method
where the $O$ term is uniform in $n\geq 1$ and $\delta>0$ small enough. Thus, for $\delta>0$ fixed and small enough, we have
@@
We are thus left to estimate the remaining two series in the right-hand side of (72). Until the end of the proof, $\kappa$ is any positive number which depends on $\delta$ and $x_{0}$, changing line to line, small enough so that the written asymptotics are valid. To bound above the first series we use a comparison between series and integrals:
@@
and we conclude, similarly as in the previous step, that